# Adversarial Robustness Limits via Scaling-Law and Human-Alignment Studies

### v1

@@ 1 Introduction
Broadly, our major contributions are as follows [^1] Appendix A provides a detailed list of our contributions and their implications, with links to relevant sections to aid navigation.:
@@ 2 Related Work
Models not trained to be robust can also be used to generate adversarial data that fool humans, though humans are only fooled by such images when given limited time to classify them (Elsayed et al., 2018). In Gaziv et al. (2023) and in this work, adversarial data is generated using robust networks, and humans may not deduce the correct category even with extra time. The ability to trick humans using images that trick robust neural networks may not be surprising given the overlap in frequencies relied upon by each system to make classifications (Subramanian et al., 2023), [^2] Humans actually use a subset of the frequencies used by neural networks. See Appendix C for more detailed discussion. the potential for a shared mechanism for robustness in humans and machines (Harrington & Deza, 2021), and the susceptibility of biological, primate neurons to adversarial attacks (Guo et al., 2022). However, we identify that such human susceptibility causes benchmarking of robustified networks (Croce et al., 2020) to produce adversarial perturbations that induce changes in human labeling: this means benchmarking does not return a fraction of “gold-standard” human performance and instead poses the robustness problem as one that is unsolvable by networks with human-level robustness. Supporting this unsolvability on CIFAR10, the human limits we find (Figure 1’s red line) match our scaling laws’ asymptotes (e.g., the FID=0 limit is the top of the shaded green area in Figure 1).
@@ 3.3.1 Compute-Efficient Adversarial Training
    A new SOTA of $73.71$% AutoAttack accuracy with a WideResNet-94-16 trained on 500M samples.
@@
    A model predicted to be compute-optimal at the prior SOTA’s training FLOPs budget (WideResNet-82-8, 79M parameters) beats the prior SOTA (WideResNet-70-16, 267M parameters) by 1% AutoAttack while being over 3×3$\times$3 × more FLOP and parameter efficient.
@@
    A model trained entirely on synthetic data outperforms the prior SOTA, reaching $71.7$% AutoAttack accuracy.
@@ 4.1 Evaluating Human Performance
Figure 9: A sampling of valid and invalid adversarial images. We categorized 727 out of 2629 adversarial images as invalid based on our human and GPT-4 study. Invalid images revealed unanticipated behaviour of adversarial attacks, adding to images instances of objects with incorrect labels (“Deceptive”) and making the labeled instance imperceptible (“Ambiguous”). The ground truth classes are below6661:cat, 2:bird, 3:deer, 4:horse, 5:deer, 6:cat, 7:dog, 8:bird, 9:bird, 10:cat, 11:airplane, 12:deer, 13:cat, 14:horse, 15:airplane. Clean images are provided in Appendix G.
@@ 5.1 Limitations and Future Directions
Our human study was limited to three users. To overcome this limitation, we plan to make our adversarial data publicly available in a quiz format, allowing the community to contribute to the estimates we made in Section 4. Finally, while we provide an updated, more optimistic view of progress towards human-level robustness on CIFAR10, our scaling laws also suggest that the costs of reaching human performance are impractically high. More efficient algorithms (Bartoldson et al., 2023) for adversarial training can address these costs and open up paths to robust models.
@@ Acknowledgements
We thank Nicholas Carlini, Maksym Andriushchenko, Vikash Sehwag, and Pin-Yu Chen for helpful comments. This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344 and LDRD Program Project No. 23-ERD-030 (LLNL-CONF-860190).
@@ Appendix A Outline of Our Contributions and Their Implications
  • We also found that human performance itself on this task is far from perfect: existing attack formulations (relying on an $\ell_{\infty}$-norm bound to ensure attacks preserve the original label) concerningly allow generation of invalid images that do not fit their original label (Figures 1, 8, and 6).

### v2

@@ 1 Introduction
Our contributions (see Appendix A) are broadly as follows:00footnotetext: See code or take online quiz via https://github.com/bbartoldson/Adversarial-Robustness-Limits.
@@ 2 Related Work
Models not trained to be robust can also be used to generate adversarial data that fool humans, though humans are only fooled by such images when given limited time to classify them (Elsayed et al., 2018). In Gaziv et al. (2023) and in this work, adversarial data is generated using robust networks, and humans may not deduce the correct category even with extra time. The ability to trick humans using images that trick robust neural networks may not be surprising given the overlap in frequencies relied upon by each system to make classifications (Subramanian et al., 2023),  [^1] Humans use a subset of NN frequencies (see Appendix C). the potential for a shared mechanism for robustness in humans and machines (Harrington & Deza, 2021), and the susceptibility of biological, primate neurons to adversarial attacks (Guo et al., 2022). However, we identify that such human susceptibility causes benchmarking of robustified networks (Croce et al., 2020) to produce adversarial perturbations that induce changes in human labeling: this means benchmarking does not return a fraction of “gold-standard” human performance and instead poses the robustness problem as one that is unsolvable by networks with human-level robustness. Supporting this unsolvability on CIFAR10, the human limits we find (Figure 1’s red line) match our scaling laws’ asymptotes (e.g., the FID=0 limit is the top of the shaded green area in Figure 1).
@@ 3.3.1 Compute-Efficient Adversarial Training
    A new SOTA of $73.71$% AutoAttack ($93.68$% clean) accuracy by training WideResNet-94-16 on 500M samples.
@@
    A model predicted to be compute-optimal at the prior SOTA’s training FLOPs budget (WideResNet-82-8, 79M parameters) beats the prior SOTA (WideResNet-70-16, 267M parameters) by 1% AutoAttack, reaching $71.59$% AutoAttack and $93.11$% clean accuracy, while being over 3×3$\times$3 × more FLOP and parameter efficient.
@@
    A model trained entirely on synthetic data outperforms the prior SOTA, attaining $71.7$% AutoAttack accuracy and $93.27$% clean accuracy.
@@ 4.1 Evaluating Human Performance
Figure 9: A sampling of valid and invalid adversarial images. We categorized 727 out of 2629 adversarial images as invalid based on our human and GPT-4 study. Invalid images revealed unanticipated behaviour of adversarial attacks, adding to images instances of objects with incorrect labels (“Deceptive”) and making the labeled instance imperceptible (“Ambiguous”). The ground truth classes are below5551:cat, 2:bird, 3:deer, 4:horse, 5:deer, 6:cat, 7:dog, 8:bird, 9:bird, 10:cat, 11:airplane, 12:deer, 13:cat, 14:horse, 15:airplane. Clean images are provided in Appendix G.
@@ 5.1 Limitations and Future Directions
Our human study was limited to three users. To address this limitation, we make a subset of this study’s adversarial data publicly available via an online quiz, and we will make more data available upon request, allowing the community to contribute to the estimates we made in Section 4. Finally, while we provide an updated, more optimistic view of progress towards human-level robustness on CIFAR10, our scaling laws also suggest that the costs of reaching human performance are impractically high. More efficient algorithms (Bartoldson et al., 2023) for adversarial training can address these costs and open up paths to robust models.
@@ Acknowledgements
We thank Nicholas Carlini, Maksym Andriushchenko, Vikash Sehwag, and Pin-Yu Chen for helpful comments. We thank Youngsoo Choi for facilitating our human evaluation. This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344 and LDRD Program Project No. 23-ERD-030 (LLNL-CONF-860190).
@@ Appendix A Outline of Our Contributions and Their Implications
  • We also found that human performance itself on this task is far from perfect: existing attack formulations (relying on an $\ell_{\infty}$-norm bound to ensure attacks preserve the original label) concerningly allow generation of invalid images that do not fit their original label (Figures 1, 8, and 5).